Express the confidence interval (197.5,409.9)(197.5,409.9) in the form of ¯x±MEx¯±ME.

¯x±ME=x¯±ME= ±±
Answer #1

Given confidence Interval (197.5, 409.9) -ME 197.5 - 0 o +ME = 409.9 - © Adding O equations eve get 25 = 197.5 +409.9 ģ= 303.